if we changed our speed limit signs to metric what would probably replace 60 mi/h
Answer #1

1 mile = 1610 m

1 h = 60*60 = 3600 s

1 mi/h = 1610/3600 = 0.44704 m/s

60 mi/h = 60*0.44704 = 26.8224 m/s